OpenStudy (anonymous):

Differentiate the function w.r.t. x from first principles. y=(x^2+3)(x-2)/(x^2) i spent more than ten minutes on this, i wonder if there is a simpler method to do this.

OpenStudy (anonymous):

i took out the common factor (x-2)

OpenStudy (anonymous):

solving for y' ??

OpenStudy (anonymous):

yes, from first principles

OpenStudy (anonymous):

limx->h {[f(x+h)-f(h)]/h}

OpenStudy (anonymous):

it's lim h->0 ??

OpenStudy (anonymous):

yes! sorry!!!! i typed that incorrectly!

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

so is there a simpler method to do this?

OpenStudy (lgbasallote):

there's never an easy way for that method....

OpenStudy (lgbasallote):

the simpler method is "DERIVATIVES"

OpenStudy (anonymous):

???

OpenStudy (lgbasallote):

if it told you to use the first principle, then you have no choice but to use it. so there's no "simpler method"

OpenStudy (anonymous):

but i have to learn to differentiate different functions from first principles

OpenStudy (lgbasallote):

because it told you which method to use

OpenStudy (lgbasallote):

does that make sense?

OpenStudy (anonymous):

yes i know i have to do this by first principles, but there are still many methods to do it.. i just don't want to expand everything. should i take out the factor (x-2) or something else?

OpenStudy (lgbasallote):

insane algebra skills

OpenStudy (lgbasallote):

that's your only way to make it simple

OpenStudy (anonymous):

........

OpenStudy (lgbasallote):

you can also expand the numerator first then divide each term by x^2

OpenStudy (lgbasallote):

only then would you use the limits

OpenStudy (lgbasallote):

\[(x^2 +3)(x-2) = x^3 - 2x^2 + 3x - 6\] \[\frac{(x^2 + 3)(x-2)}{x^2} \implies \frac{x^3 - 2x^2 + 3x - 6}{x^2} \implies x - 2 + \frac 2x - \frac 6{x^2}\]

OpenStudy (lgbasallote):

i don't know if that's easier for you though

OpenStudy (anonymous):

I think so :P, it's insane

OpenStudy (anonymous):

ok .... thanks

OpenStudy (lgbasallote):

"insane algebra skills" wasn't an exaggeration. I really was serious that that is the only way to make it simpler